### v1

# External Incremental Delaunay Triangulation

Abstract
This paper introduces a Delaunay triangulation algorithm based on the external incremental method. Unlike traditional random incremental methods, this approach uses convex hull and points as basic operational units instead of triangles. Since each newly added point is outside the convex hull, there is no need to search for which triangle contains the point, simplifying the algorithm implementation. This paper details the algorithm’s data structures, implementation details, correctness proof, and comparison with other methods.
1 Introduction
Delaunay triangulation is an important problem in computational geometry, with wide applications in computer graphics, geographic information systems, and other fields. Traditional random incremental methods use triangles as basic operational units, requiring the search for which triangle contains each newly added point. The external incremental method proposed in this paper avoids this search process by using convex hull as the basic operational unit, simplifying the algorithm implementation.
2 Algorithm Overview
The basic idea of the external incremental method is: each newly added point is outside the current convex hull, therefore it cannot be inside any existing triangle. This property eliminates the need to search for which triangle contains the new point, simplifying the algorithm implementation.
3 Data Structures
3.1 Array P
$P$ is a two-dimensional array used to store a set of two-dimensional points. It satisfies $x_{i}\leq x_{j}$ when $i<j$, and if $x_{i}=x_{j}$, $i<j$, then $y_{i}\leq y_{j}$.
3.2 Index List L_ch_i
$L\_\text{ch}\_i$ is an index list whose elements point to points in $P$ that form a convex hull. It can be represented as: $L\_ch\_i=[i_{1},i_{2},\ldots,i_{m}]$. Here, $i_{k}$ is the index of points in $P$ that form the convex hull.
3.3 Dictionary D_tri
D_tri is a dictionary representing the Delaunay triangulation. Its keys are pairs of point indices, and its values are lists of indices of the third points that form triangles with the key pair. Each list contains one or two integer values, which are indices of points that form triangles with the first two points. The specific form is: $\{(i_{1},i_{2})\mapsto[i_{3},i_{4}],\cdots,(i_{5},i_{6})\mapsto[i_{7}],\cdots\}$.
This indicates that in $P$, points with indices $i_{1},i_{2},i_{3}$ form a triangle, and points with indices $i_{1},i_{2},i_{4}$ also form a triangle. Specifically, the line segment $P[i_{1}]P[i_{2}]$ is an edge of the triangle, while points $P[i_{3}]$ and $P[i_{4}]$ are vertices opposite to this edge.
4 Algorithm Implementation
4.1 Main Algorithm
Data: Sorted point set P in 2D plane
Result: Convex hull L of this point set, Delaunay triangulation
$i\leftarrow 2$
repeat
       $L\_ch\_i,D\_tri,i\_xmax\leftarrow\text{Try to create a triangle with P[0,1,...,i]}(P,i)$
      $i\leftarrow i+1$
until L is not empty;
while $i<len(P)$ do
       $L,i\_xmax\leftarrow\text{Add point i to current Delaunay triangulation}(P,L\_ch\_i,D\_tri,i\_xmax,i)$;
       $i\leftarrow i+1$;
end while
return $L$;
Algorithm 1 External Incremental Method
Data: P,A,D,a1,i_p
Result: L, i_xmax, new Delaunay triangulation (stored in D_tri, not explicitly returned)
Find upper tangent point a1 and lower tangent point a2 between P[i] and current convex hull
Initialize N as empty list %N acts as a stack
Push indices of all points between a1 and a2 into N
L_i=[A[a1]] %L_i records exposed points
find_border_while(P,D,N,L_i,A[a1],i_p) %Implement erosion using while loop
Connect exposed points with p to update L, D_tri
Algorithm 2 Add point i to current Delaunay triangulation
Input: Point set $P$, dictionary $D$, list $N$, list $L\_i$, index $i\_b$, index $i\_p$
Output: Updated $D$, $N$, $L\_i$
while $N$ is not empty do
       if $(i\_b,N[-1])$ is an edge then
             $i\_d\leftarrow D[i\_b,N[-1]][0]$;
             if The point to be added is in its determined triangle then
                   Delete this edge
                  N.append(i_d)
                  continue
             end if
       end if
[…]
      L_i.append(i_b)
end while
Algorithm 3 find_border_while function
4.2 Algorithm Illustration
(a) Find upper and lower tangent points of convex hull
(b) Determine if the new point is inside the circumcircle of the triangle
formed by the line segment between tangent points
(c) If an edge is deleted, two new edges need to be checked
(d) Connect exposed points with new point
A new convex hull is formed, proceed to next iteration
Figure 1: Sorted incremental method for Delaunay triangulation
5 Point in Circumcircle Test
Notation: Let A be the point to be added, B and C be the points to be processed. D is the point corresponding to line segment BC, $\triangle BCD$ is a triangle. $\angle{A}=\angle{BAC}$, $\angle{D}=\angle{CDB}$
Note that point A to be added must be outside the convex hull. Therefore, it must be on the opposite side of point D corresponding to the line segment BC being processed. Also, note that neither $\angle A$ nor $\angle D$ is greater than $\pi$. With these two conditions, we can determine if point p is inside the circumcircle of $\triangle BCD$ by checking if $\angle{A}+\angle{D}$ is greater than $\pi$.
$\angle{A}+\angle{D}$ $>\pi$ (1)
$\angle{A}$ $>\pi-\angle{D}$ (2)
$\cos A$ $<-\cos D$ (3)
$\cos A+\cos D$ $<0$ (4)
$\frac{\textbf{AB}\cdot\textbf{AC}}{|AB||AC|}+\frac{\textbf{BD}\cdot\textbf{CD}}{|BD||CD|}$ $<0$ (5)
$|\textbf{AB}\cdot\textbf{AC}+|AB||AC|\textbf{BD}\cdot\textbf{CD}$ | BD ⋅ CD $<0$ (6)
Proposition 1.
For $\triangle BCD$, if A and D are on opposite sides of BC, then if $\angle{A}+\angle{D}>\pi$, A is inside the circumcircle of BCD. If $\angle{A}+\angle{D}<\pi$, A is outside the circumcircle of BCD.
(a) A is inside the circumcircle of BCD
(b) A is outside the circumcircle of BCD
Figure 2: A and D are on opposite sides of BC
Proof.
Let O be the circumcenter of $\triangle BCD$.
If A is inside the circumcircle of BCD, extend CA to intersect the circle at point E. $\angle{BEC}+\angle{BDC}=\pi$. $\angle{BAC}=\angle{BEC}+\angle{EBA}>\angle{BEC}$. Therefore $\angle{BAC}+\angle{BDC}>\pi$.
If A is outside the circumcircle of BCD, connect OA to intersect the circle at point F. $\angle{BFC}+\angle{BDC}=\pi$.
$\angle{BFC}=\angle{BFO}+\angle{OFC}$, $\angle{BAC}=\angle{BAO}+\angle{OAC}$.
$\angle{BFO}=\angle{BAO}+\angle{ABF}>\angle{BAO}$, $\angle{OFC}=\angle{OAC}+\angle{ACF}>\angle{OAC}$.
Therefore $\angle{BAC}<\angle{BFC}$, $\angle{A}+\angle{D}<\pi$. ∎
Note that for $\triangle BCD$, when A and D are on the same side of BC, the above does not hold. The question of whether A is inside the circumcircle of BCD can be solved by computing the Cayley-Menger determinant, but this is a $5*5$ determinant. Clearly, Algorithm 1 ensures that A and D are always on opposite sides of BC at each step, so we can use inequality 6. Generally, division is more prone to precision loss than multiplication, and some processor instruction sets (like some ARM architectures) don’t even include division operations, so inequality 6 is converted to a form using only multiplication.
[…]
6 Correctness Proof
This can be clearly seen through the dual graph of Delaunay triangulation, the Voronoi diagram.
Theorem 1.
In Algorithm 1, inequality 6 always holds. That is, point A to be added is always on the opposite side of point D corresponding to line segment BC being processed.
Theorem 2.
Through a series of edge flips, any two triangulations of the same planar point set can be transformed into each other.[1]
The Delaunay graph is dual to the Voronoi diagram. Each Delaunay triangle corresponds to a Voronoi region. Edges in the Delaunay graph correspond to edges in the Voronoi diagram. Points in the Delaunay graph correspond to sites in the Voronoi diagram. By the property of Voronoi diagrams: a site $p_{i}$’s Voronoi region $\text{VR}(p_{i})$ is open if and only if $p_{i}$ is on the convex hull of point set $P$. Therefore, in Algorithm 1, after adding each new point, some Voronoi sites that were previously on the convex hull become interior points. Their Voronoi regions become closed. By $\text{VR}(p_{i})=\{q|~{}|qp_{i}|\leq|qp_{j}|,~{}\forall j\neq i,~{}p_{i},p_{j}\in P\}$[3, page 51], $\text{VR}(p_{i})=\bigcap\limits_{i\neq j}h(p_{i},p_{j})$, $h(p_{i},p_{j})=\{q|~{}|qp_{i}|\leq|qp_{i}|\}$. That is, changes to a site’s Voronoi region are caused by the newly added point. Therefore, there must exist a Voronoi edge whose associated sites are the original open Voronoi region’s site and the newly added site. That is, new edges in the Delaunay graph must be connected to the newly added point.
7 Analysis and Comparison
The point addition operation in this method is similar to the random incremental method[2]. However, compared to the random incremental method, this method does not use a fictitious large outer triangle, thus eliminating a series of operations and assumptions related to this fictitious triangle. Additionally, this method does not involve determining which triangle contains a point. Therefore, there is no need to dynamically maintain a triangle search tree during construction.